the vertex of this parabola is at 5,5 the x value is 6 and the y is -1. WHat is the coefficient of the squared expression in the parabola's equation?
Answer by lwsshak3(11628)

the vertex of this parabola is at 5,5 the x value is 6 and the y is -1. WHat is the coefficient of the squared expression in the parabola's equation?
***
Vertex form of equation for a parabola: y=A(x-h)^2+k, (h,k)=(x,y) coordinates of the vertex. A is a coefficient that affects the slope or steepness of the curve.
For given problem:
-1=A(6-5)^2+5
-1=A(1)^2+5
A=-6
Equation: y=-6(x-5)^2+5
This is a parabola that opens downward with vertex at (5,5)